# A Tutorial on MIMO-OFDM ISAC: From Far-Field to Near-Field

### v1

@@ II-A1 Primary ISAC Tasks and KPIs
To evaluate the performance of ISAC systems, several KPIs have been specified for C&S services. The performance of communication tasks can be assessed from two key perspectives: efficiency and reliability. Efficiency is used to assess the amount of information successfully delivered from the transmitter to the receiver given limited resources. Reliability refers to the ability of the system to withstand adverse communication environments, such as channel fading, interference, and noise. This metric focuses on transmission quality, aiming to ensure that information is not corrupted or lost during transmission. On the other hand, sensing tasks can be generally classified into three categories, i.e., target detection, parameter estimation, and object imaging, with different KPIs. Target detection refers to determining the presence or absence of targets obscured by clutter and interference [43]. Parameter estimation refers to the problem of extracting relevant parameters from target reflections. For unbiased estimators, the Cramér-Rao bound (CRB) characterizes the theoretical lower bound on mean squared error (MSE) between the actual parameters and their estimates [22, 40]. The number of sensing degrees of freedom (SDoF) is also an important KPI for parameter estimation, characterizing the maximum number of targets whose parameters can be estimated. For multi-target sensing, resolution is an essential KPI that refers to the ability of the system to distinguish objects with similar parameter values. Object imaging is a sensing task that aims to determine the type and shape of sensed objects. Imaging deviation is used as a KPI in this case to measure the accuracy of the imaging process [44]. In addition, image entropy and image contrast can be employed to evaluate the quality of the imaging results [45].
@@ II-A2 ISAC transmitter
The design of ISAC transmission systems involves key considerations aimed at optimizing both C&S functions, such as waveform design, resource allocation, antenna configuration, and beam management. Communication is typically BS-centric with single-hop propagation, whereas sensing is governed by the classical radar distance equation [46]. A dedicated waveform that can achieve higher degrees of integration between C&S without interference is crucial. One straightforward solution is to schedule C&S on orthogonal resource blocks, so that they do not interfere with each other [47, 48, 49]. However, orthogonal resource allocation usually suffers from poor SE and EE. Therefore, in order to maximize the integration gain between C&S, it is more favorable to design a fully unified ISAC waveform.
Multi-antenna technology has evolved from MIMO in fourth-generation (4G) networks to massive MIMO in fifth-generation (5G) networks. For future 6G systems, several emerging technologies may greatly improve the number of available spatial DoFs, such as extremely large-scale MIMO (XL-MIMO) [50, 51], sparse MIMO [52, 53], modular MIMO [54, 55] and fluid [56]/movable antennas [57, 58, 59, 60]. As array apertures increase, the beams of future XL-MIMO and mmWave/THz ISAC systems may be extremely narrow, increasing the challenges of beam management in achieving high gain and avoiding misalignment. In the literature, transmit beamformers are often designed to minimize the discrepancy between the resulting and targeted beampattern for sensing, while minimizing multi-user interference [61, 62, 63, 64]. To further address the complexity and cost incurred by fully digital MIMO, hybrid beamforming has been studied for enhancing ISAC performance in full-duplex systems [65, 66].
@@ II-A3 ISAC Channel
The mathematical models for C&S channels differ significantly. Communication channel models are typically stochastic, focusing on signal transmission quality, while sensing channels are deterministic, incorporating target scattering. The transmitters and receivers in communication systems are usually assumed to be spatially separated, while the sensing transmitters and receivers can be collocated or separated, corresponding to “monostatic sensing” or “bistatic sensing”, respectively. While multipath environments can enhance the spatial diversity of communication systems, radar sensing usually requires a line-of-sight (LoS) link for target detection. Channel fading models are also different for C&S. For instance, large- and small-scale fading are typically used to model the propagation effects in communication channels, while for radar sensing, channel fading is usually characterized in terms of target scattering characteristics, such as the radar cross section (RCS).
@@
With the ever-increasing array apertures and the use of mmWave/terahertz frequencies, the (radiative) near-field region will significantly expand, even up to several hundreds of meters. Thus, it is crucial to consider algorithm development and performance for ISAC systems that operate in near-field scenarios. In the near-field region, the more accurate USW model needs to be considered to characterize both the phase and amplitude variations across array elements [51, 50]. The USW models for ISAC enable the joint estimation of target angle and range even with only single-anchor array processing.
Furthermore, as the operating frequency increases, ISAC signals become more susceptible to blockages, which can significantly degrade C&S performance. Fortunately, several techniques have been proposed to address this issue, such as using intelligent reflecting surfaces (IRSs) [70, 71] and fully passive metal reflectors [72] that establish virtual wireless links. Furthermore, one can also exploit prior knowledge of the local environment to enhance C&S performance. For example, approaches based on a channel channel knowledge map (CKM) have been proposed to acquire CSI for achieving environment-aware C&S [73, 74].
@@ II-A4 ISAC receiver
Advanced signal processing techniques need to be developed for the ISAC receiver to achieve favorable C&S performance, addressing issues such as transceiver synchronization, interference management, clutter suppression and target parameter estimation. For monostatic sensing, the receiver is collocated with the ISAC transmitter, and thus it has full knowledge of the transmitted signals. As such, the sensing receiver does not need to perform a dedicated synchronization or demodulation of the communication symbols. However, self-interference occurs due to signal leakage between the transmitter and receiver, which may lead to non-linear distortion and difficulty in sensing weak targets [75, 76]. For bistatic sensing, the ISAC receiver needs to estimate and compensate for clock bias, and the communication symbols need to be demodulated before performing sensing.
Typical parameters that need to be estimated include the angle of arrival (AoA), distance, and velocity of targets. Sensing receivers usually estimate the distance and velocity parameters by analyzing the propagation delays and Doppler shifts of the echo signals, respectively. For AoA estimation, multiple antennas are needed to provide spatial resolution. These estimation problems have been studied for many decades. For example, data-based algorithms such as Capon beamforming [77] and amplitude and phase estimation (APES) [78] can be used to improve the angle estimation accuracy by optimizing the beamforming gain at the receiver. Subspace-based techniques, such as multiple signal classification (MUSIC) [79] and estimation of signal parameters via rotational invariance techniques (ESPRIT) [80], exploit the orthogonality between the signal subspace and the noise subspace to achieve super-resolution angle estimation. Compared to subspace-based algorithms, CS and sparse recovery methods exploit channel sparsity to reconstruct the transmitted signals and extract target parameters. However, sparse recovery inherently involves NP-hard $\ell_{0}$-norm optimization, which is typically approximated using greedy algorithms like orthogonal matching pursuit (OMP) [81]. In [82], the authors demonstrated that the $\ell_{0}$-norm problem can be relaxed to an $\ell_{1}$-norm formulation to achieve tractable solutions involving convex optimization algorithms based on the $\ell_{1}$-norm [83, 84] or the $\ell_{1}$-singular value decomposition (SVD) [85]. While both $\ell_{0}-$ and $\ell_{1}-$norm minimization require discrete grids that introduce potential basis mismatch and degraded sensing accuracy, atomic norm minimization (ANM) [86, 87] avoids grid dependency but incurs higher computational complexity. ISAC techniques that use super-resolution sensing algorithms fall under the heading integrated super-resolution sensing and communication ($\rm{IS^{2}AC}$) [25, 88, 89], and can provide highly accurate parameter estimates. However, such algorithms are known to be sensitive to array modeling errors, and they are costly to implement for large-scale arrays. These factors must be taken into account when considering them for ISAC scenarios.
@@ II-B Potential ISAC Waveforms and OFDM ISAC
To meet the diverse C&S requirements of 6G ISAC applications, using an appropriate waveform is critical. For high communication performance, such as high data rates and low BER, waveforms should feature high SE, low peak-to-average power ratio (PAPR), robustness to Doppler effects and phase noise, resilience to nonlinear power amplifier noise, low signal processing complexity, and compatibility with legacy MIMO systems. For radar sensing, waveforms should enhance resolution, estimation accuracy, clutter rejection, and interference resistance by exhibiting low PAPR and favorable ambiguity function characteristics, such as narrow mainlobe width, low peak-to-sidelobe ratio (PSLR) or low integrated sidelobe ratio (ISLR). However, achieving all these attributes simultaneously is challenging, resulting in trade-offs between C&S performance. Based on these trade-offs, one can classify waveform design approaches as either sensing-centric, communication-centric, or joint designs.
TABLE I: Comparison of advantages and limitations for various ISAC waveforms.
@@
In this paper, we focus on communication-centric ISAC systems. Communication-centric approaches leverage existing communication waveforms with various modifications to enable or enhance their sensing capability. However, the inherent randomness of communication waveforms typically degrades sensing performance due to issues like high PAPR, random autocorrelation properties, and reduced resolution [16]. Communication waveforms can be categorized as either single-carrier or multi-carrier. Single-carrier waveforms feature low PAPR and minimal nonlinear distortion [104], making them suitable for scenarios with limited coverage but high communication quality requirements. Examples include single-carrier frequency division multiple access (SC-FDMA) [90] methods, such as DFT-spread-OFDM (DFT-s-OFDM) [16, 91], widely used in the uplink of 4G and 5G systems. The authors of [23] and [92] studied single-carrier ISAC using the delay-Doppler alignment modulation (DDAM) technique, which eliminates inter-symbol interference (ISI) through path-based beamforming and delay compensation. On the other hand, multi-carrier waveforms enable increased communication rates and SE [104]. OFDM [16, 25] is a quintessential example of multi-carrier waveforms and is ubiquitously used in 4G and 5G systems, featuring a “thumbtack-shaped” ambiguity function that provides an inherent sensing capability. Additionally, emerging multi-carrier waveforms like orthogonal time frequency space (OTFS) [26], orthogonal chirp division multiplexing (OCDM) [27], and affine frequency division multiplexing (AFDM) [28] exhibit greater robustness against Doppler for high-mobility ISAC applications at the cost of higher implementation complexity.
@@
Table I presents the main types of ISAC waveforms along with their advantages and disadvantages. In summary, sensing-centric waveforms are limited by low communication rates. For joint designs, multiplexing methods suffer from inefficient radio resource utilization, while fusion methods entail high implementation complexity and/or complex optimization problems. For communication-centric waveforms, emerging technologies like DDAM, OTFS, OCDM, and AFDM exhibit superior sensing performance in high-mobility scenarios but are often accompanied by complex signal processing and high computational demands. Consequently, OFDM remains a highly promising waveform for future ISAC wireless networks. For MIMO-OFDM ISAC systems, the use of XL-MIMO and large time-frequency bandwidths significantly enhances the sensing performance, while they simultaneously introduce substantial computational complexity for parameter estimation. This challenge becomes particularly pronounced when performing joint estimation across the spatial, delay, and Doppler domains, compounded by the additional complexity of near-field parameter estimation. Thus, in this paper, we present a tutorial overview of rigorous and unified models of MIMO-OFDM ISAC, the methods and corresponding algorithms for parameter estimation across the three domains, as well as their performance and complexity. Furthermore, the extension of far-field methods to near-field scenarios will be elaborated.
@@ III Fundamental ISAC Models
Considering the universal applicability of ISAC frameworks to all possible waveforms, we first introduce various ISAC modes, characterizing the relationships among transmitter and receiver signals and considering the required signal processing. As shown in Fig. 1, 3GPP standardization bodies have discussed six basic ISAC modes [105], which can be divided into two categories depending on whether the sensing signal is received and processed at the BS or UE side. For BS-side sensing, as shown in Fig. 1 (a)-(c), ISAC signals could be sent by either the BS itself, a UE, or another BS, corresponding to the BS monostatic, UE-to-BS bistatic, and BS-to-BS bistatic modes, respectively. Similarly, for UE-side sensing, as shown in Fig. 1 (d)-(f), ISAC signals could be sent by either the UE itself, a BS, or another UE, corresponding to the UE monostatic, BS-to-UE bistatic, and UE-to-UE bistatic modes, respectively. In the following, we present the signal input-output relationships for C&S based on a unified ISAC model that is applicable to all the six modes above.
@@ III-A Unified ISAC Model
Fig. 2 shows a generic ISAC setup that includes the aforementioned six modes as special cases. It consists of three components, namely an ISAC transmitter (ISAC-TX), a communication receiver (Com-RX) and a sensing receiver (Sen-RX). The ISAC-TX transmits ISAC signals to the Com-RX to provide communication services, and to the Sen-RX for sensing $K$ targets whose locations denoted as $\mathbf{q}_{k},k=1,2,\cdots,K$. Denote the number of antennas at the ISAC-TX, Com-RX, and Sen-RX by $M_{t}$, $M_{c}$, and $M_{r}$, respectively. For sensing, we assume that each target has an LoS path with the ISAC-TX and the Sen-RX, while such an assumption is not needed for communication. The array response vector between the ISAC-TX and target $k$ is denoted as $\mathbf{a}_{\mathrm{T}}(\mathbf{q}_{k})\in\mathbb{C}^{M_{t}\times 1}$, which is a function of the sensing target location $\mathbf{q}_{k}$ in general. Similarly, the RX array response vector is denoted as $\mathbf{a}_{\mathrm{R}}(\mathbf{q}_{k})\in\mathbb{C}^{M_{r}\times 1}$. Specific expressions for $\mathbf{a}_{\mathrm{T}}(\mathbf{q}_{k})$ and $\mathbf{a}_{\mathrm{R}}(\mathbf{q}_{k})$ in far-field and near-field scenarios will be provided in Section III-B.
@@
where $\ast$ denotes convolution, and $\mathbf{n}\bigl{(}t\bigr{)}\sim CN\bigl{(}0,\mathbf{\sigma}^{2}\mathbf{I}_{M_{c}}\bigr{)}$ is circularly symmetric additive white Gaussian noise (AWGN) with zero mean and variance $\sigma^{2}$. Communication signal processing methods remain consistent with traditional communication systems and will not be detailed here.
For target sensing, the ISAC signal sent by the ISAC-TX is reflected/scattered by the $K$ targets and other scatterers in the surrounding environment, and then received by the Sen-RX. An LoS link between the ISAC-TX and Sen-RX may also exist. Thus, the sensing signal $\mathbf{y}_{s}\left(t\right)\in\mathbb{C}^{M_{r}\times 1}$ received by the $M_{r}$-antenna Sen-RX array can be expressed as
@@
×$\times e^{j2\pi\left(f_{\Delta}+\upsilon_{k}\right)t}+\mathbf{H}_{\mathrm{Clutter}}\left(t,\tau\right)\ast\mathbf{x}\left(t\right)+\tilde{\mathbf{z}}\left(t\right),$
where $\alpha_{k}$, $\tau_{k}$, and $\upsilon_{k}$ denote the complex-valued channel coefficient, the propagation delay, and Doppler frequency of the path corresponding to the $k$-th sensing target, respectively, $\tau_{\Delta}$ and $f_{\Delta}$ are the symbol time offset (STO) and carrier frequency offset (CFO) caused by the difference between the reference clock and local oscillators (LO) at the ISAC-TX and Sen-RX. Note that we may assume $\tau_{\Delta}=0$ and $f_{\Delta}=0$ for monostaic ISAC modes, while $\tau_{\Delta}\neq 0$ and $f_{\Delta}\neq 0$ in general for bistatic ISAC modes. The term $\mathbf{H}_{\mathrm{Clutter}}\left(t,\tau\right)$ is the channel impulse response of the clutter, which consists of the signals scattered/reflected by other scatterers irrelevant to the $K$ sensing targets and the LoS link signal between the ISAC-TX and Sen-RX. Note that for monostatic ISAC systems, the ISAC-TX and Sen-RX are collocated and operate in full-duplex mode to achieve maximum SE, making them vulnerable to self-interference (SI) due to imperfect isolation. The residual SI after the application of various cancellation techniques can also be considered as a component of the clutter. Finally, $\tilde{\mathbf{z}}\left(t\right)$ is the AWGN at the Sen-RX.
Note that except for some differences in the number of antennas and computational capabilities, BS-side and UE-side sensing are symmetric in terms of signal modeling and processing. Therefore, in the following, we will present detailed models and signal processing techniques for ISAC systems of the three BS-side sensing modes based on the unified ISAC model, while similar results can be applied to UE-side sensing as well. To this end, we first show that the generic ISAC setup in Fig. 2 and the resulting signal in (3) can be reduced to the three BS-side sensing modes shown in Figs. 1 (a)-(c).
@@
TABLE II: The unified ISAC model and its special cases corresponding to the six basic modes.
@@
  • BS monostatic ISAC mode: For BS monostatic ISAC mode, the ISAC-TX and Sen-RX shown in Fig. 2 are collocated at the BS. In order to maximize the SE, the ISAC BS needs to employ in-band full-duplex operation for simultaneous downlink ISAC signal transmission and uplink sensing signal reception. Note that while collocated at the BS, we assume that the ISAC-TX and Sen-RX use different sets of antennas for ease of SI cancellation. The Com-RX in Fig. 2 corresponds to a UE. In this case, the generic ISAC setup in Fig. 2 reduces to the BS monostatic setup in Fig. 3(a). Furthermore, the ISAC signal $\mathbf{x}(t)$ in (1)-(3) is transmitted by the BS, and is denoted as $\mathbf{x}_{\mathrm{BS}}(t)\in\mathbb{C}^{M_{t}\times 1}$.

### v2

@@ II-A1 Primary ISAC Tasks and KPIs
To evaluate the performance of ISAC systems, several KPIs have been specified for C&S services. The performance of communication tasks can be assessed from two key perspectives: efficiency and reliability. Efficiency is used to assess the amount of information successfully delivered from the transmitter to the receiver given limited resources. Reliability refers to the ability of the system to withstand adverse communication environments, such as channel fading, interference, and noise. This metric focuses on signal quality, aiming to ensure that information is not corrupted or lost during transmission. On the other hand, sensing tasks can be generally classified into three categories, i.e., target detection, parameter estimation, and object imaging, with different KPIs. Target detection refers to determining the presence or absence of targets obscured by clutter and interference [43]. Parameter estimation refers to the problem of extracting relevant parameters from target reflections. For unbiased estimators, the Cramér-Rao bound (CRB) characterizes the theoretical lower bound on mean squared error (MSE) between the actual parameters and their estimates [22, 40]. The number of sensing degrees of freedom (SDoF) is also an important KPI for parameter estimation, characterizing the maximum number of targets whose parameters can be estimated. For multi-target sensing, resolution is an essential KPI that refers to the ability of the system to distinguish objects with similar parameter values. Object imaging is a sensing task that aims to determine the type and shape of sensed objects. Imaging deviation is used as a KPI in this case to measure the accuracy of the imaging process [44]. In addition, image entropy and image contrast can be employed to evaluate the quality of the imaging results [45].
@@ II-A2 ISAC Transmitter
The ISAC transmitter design involves optimizing both C&S functions, including waveform design, resource allocation, antenna configuration, and beam management. Communication is typically based on the single-hop propagation between BS and user equipment (UE), whereas sensing is governed by the classical radar equation [46]. A dedicated waveform that can achieve a higher degree of integration between C&S without interference is crucial. One straightforward solution is to schedule C&S on orthogonal resource blocks, so that they do not interfere with each other [47, 48, 49]. However, orthogonal resource allocation usually suffers from poor SE and EE. Therefore, in order to maximize the integration gain between C&S, it is more favorable to design a fully unified ISAC waveform.
Multi-antenna technology has evolved from MIMO in fourth-generation (4G) networks to massive MIMO in fifth-generation (5G) networks. For future 6G systems, several emerging technologies may further increase the number of available spatial DoFs, such as extremely large-scale MIMO (XL-MIMO) [50, 51], sparse MIMO [52, 53], modular MIMO [54, 55], and fluid [56]/movable antennas [57, 58, 59, 60]. Millimeter-wave (mmWave) and terahertz (THz) frequency bands offer abundant spectral resources, where the large available bandwidth empowers both high-rate communications and high-resolution delay/distance sensing. Furthermore, the shorter wavelength allows for the deployment of more antenna elements within the same physical aperture, thereby achieving higher MIMO spatial multiplexing and beamforming gains. As array apertures increase, the beams of future XL-MIMO and mmWave/THz ISAC systems may be extremely narrow, increasing the challenges of beam management in achieving high gain and avoiding misalignment. In the literature, transmit beamformers are often designed to minimize the discrepancy between the resulting and targeted beampattern for sensing, while minimizing multi-user interference [61, 62, 63, 64]. To further address the complexity and cost incurred by fully digital MIMO, hybrid beamforming has been studied for enhancing ISAC performance in full-duplex systems [65, 66].
@@ II-A3 ISAC Channel
The mathematical models for C&S channels differ significantly. Communication channel models are typically stochastic, focusing on signal transmission quality, while sensing channels are deterministic, incorporating target scattering. The transmitters and receivers in communication systems are usually assumed to be spatially separated, while the sensing transmitters and receivers can be collocated or separated, corresponding to “monostatic sensing” or “bistatic sensing”, respectively. While multipath environments can enhance the spatial diversity of communication systems, radar sensing usually requires a line-of-sight (LoS) link for target detection. Channel fading models are also different for C&S. For instance, large- and small-scale fading are typically used to model the propagation effects in communication channels, while for radar sensing, signal propagation is also characterized in terms of target scattering characteristics, such as the radar cross section (RCS).
@@
With ever-increasing array apertures and the use of mmWave/THz frequencies, the (radiative) near-field region will significantly expand, even up to several hundreds of meters. Thus, it is crucial to consider algorithm development and performance for ISAC systems that operate in near-field scenarios. In the near-field region, the more accurate USW model needs to be considered to characterize both the phase and amplitude variations across array elements [51, 50]. The USW models for ISAC enable joint estimation of target angle and distance even with only spatial-domain array signal processing, thus reducing the requirements of high-precision synchronization.
Furthermore, as the operating frequency increases, ISAC signals become more susceptible to blockages, which can significantly degrade C&S performance. Fortunately, several techniques have been proposed to address this issue, such as using intelligent reflecting surfaces (IRSs) [70, 71] and fully passive metal reflectors [72] that establish virtual wireless links. Furthermore, one can also exploit prior knowledge of the local environment to enhance C&S performance. For example, approaches based on a channel knowledge map (CKM) have been proposed to acquire CSI for achieving environment-aware C&S [73, 74].
@@ II-A4 ISAC Receiver
Advanced signal processing techniques need to be developed for the ISAC receiver to achieve favorable C&S performance, addressing issues such as transceiver synchronization, interference management, clutter suppression, and target parameter estimation. For monostatic sensing, the receiver is collocated with the ISAC transmitter, and thus it has full knowledge of the transmitted signals. As such, the sensing receiver does not need to perform a dedicated synchronization or demodulation of the communication symbols. However, self-interference occurs due to signal leakage between the transmitter and receiver, which may lead to non-linear distortion and difficulty in sensing weak targets [75, 76]. For bistatic sensing, the ISAC receiver needs to estimate and compensate for clock bias, and the communication symbols need to be demodulated before performing sensing.
Typical parameters that are estimated include the angle of arrival (AoA), distance, and velocity of targets. Sensing receivers usually estimate the distance and velocity parameters by analyzing the propagation delays and Doppler shifts of the echo signals, respectively. For AoA estimation, multiple antennas are needed to provide spatial resolution. These estimation problems have been studied for many decades. For example, data-based algorithms such as Capon beamforming [77] and amplitude and phase estimation (APES) [78] can be used to improve the angle estimation accuracy by optimizing the beamforming gain at the receiver. Subspace-based techniques, such as multiple signal classification (MUSIC) [79] and estimation of signal parameters via rotational invariance techniques (ESPRIT) [80], exploit the orthogonality between the signal subspace and the noise subspace to achieve super-resolution angle estimation. Compared to subspace-based algorithms, CS and sparse recovery methods exploit channel sparsity to reconstruct the transmitted signals and extract target parameters. However, sparse recovery inherently involves an NP-hard $\ell_{0}$-norm optimization, which is typically approximated using greedy algorithms like orthogonal matching pursuit (OMP) [81]. ISAC techniques that use super-resolution sensing algorithms fall under the heading integrated super-resolution sensing and communication ($\rm{IS^{2}AC}$) [25, 82, 83], and can provide highly accurate parameter estimates. However, such algorithms are known to be sensitive to array modeling errors, and they are costly to implement for large-scale arrays. These factors must be taken into account when considering them for ISAC scenarios.
@@ II-B Potential ISAC Waveforms and OFDM ISAC
To meet the diverse C&S requirements of 6G ISAC applications, using an appropriate waveform is critical. For high communication performance with high data rates and low BER, waveforms should feature high SE, low peak-to-average power ratio (PAPR), robustness to Doppler effects and phase noise, resilience to nonlinear power amplifier noise, low signal processing complexity, and compatibility with legacy MIMO systems. For radar sensing, waveforms should enhance resolution, estimation accuracy, clutter rejection, and interference resistance by exhibiting low PAPR and favorable ambiguity function characteristics, such as narrow mainlobe width, low peak-to-sidelobe ratio (PSLR) or low integrated sidelobe ratio (ISLR). However, achieving all these attributes simultaneously is challenging, resulting in trade-offs between C&S performance. Based on these trade-offs, one can classify waveform design approaches as either sensing-centric, communication-centric, or joint designs.
TABLE III: Comparison of advantages and limitations for various ISAC waveforms.
@@
In this paper, we focus on communication-centric ISAC systems. Communication-centric approaches leverage existing communication waveforms with various modifications to enable or enhance their sensing capability. However, the inherent randomness of communication waveforms typically degrades sensing performance due to issues like high PAPR, random autocorrelation properties, and reduced resolution [16]. Communication waveforms can be categorized as either single-carrier or multi-carrier. Single-carrier waveforms feature low PAPR and minimal nonlinear distortion [98], making them suitable for scenarios with limited coverage but high communication quality requirements. Examples include single-carrier frequency division multiple access (SC-FDMA) [84] methods, such as DFT-spread-OFDM (DFT-s-OFDM) [16, 85], widely used in the uplink of 4G and 5G systems. The authors of [23] and [86] studied single-carrier ISAC using delay-Doppler alignment modulation (DDAM), which eliminates inter-symbol interference (ISI) through path-based beamforming and delay compensation. On the other hand, multi-carrier waveforms enable increased communication rates and SE [98]. OFDM [16, 25] is a quintessential example of multi-carrier waveforms and is ubiquitously used in 4G and 5G systems, featuring a “thumbtack-shaped” ambiguity function that provides an inherent sensing capability. However, OFDM also suffers from inherent drawbacks such as high PAPR and sensitivity to Doppler spread, which result in increased hardware cost and degraded C&S performance in high-mobility scenarios.
Several emerging multi-carrier waveforms, such as orthogonal time frequency space (OTFS) [26], orthogonal chirp division multiplexing (OCDM) [27], and affine frequency division multiplexing (AFDM) [28], have been proposed to overcome the limitations of OFDM. Specifically, OTFS maps communication symbols to the delay-Doppler domain and extends them over the entire time-frequency domain via a two-dimensional (2D) transform. This approach fully exploits the sparsity of wireless channels, achieving high diversity gain and demonstrating more robust C&S performance in high-mobility scenarios. Based on the Fourier orthogonal basis of OFDM, OCDM and AFDM introduce orthogonal chirp and affine chirp-like phase rotations, respectively, which shift signal orthogonality from the frequency domain to the joint time-frequency domain, thereby enhancing robustness against multipath fading and Doppler spread. Compared to OTFS and OCDM, AFDM can flexibly adjust two affine parameters to match time-frequency selective channels, ensuring orthogonality and full diversity gain, thus achieving better C&S performance. However, these waveforms introduce additional 2D transforms, resulting in more complex transceiver architectures and increased computational complexity.
@@
Table IIIpresents the main types of ISAC waveforms along with their advantages and disadvantages. In summary, sensing-centric waveforms are limited by low communication rates. For joint designs, multiplexing methods, such as time division multiplexing (TDM) [3], frequency division multiplexing (FDM) [88], code division multiplexing (CDM) [89], and space division multiplexing (SDM) [90], suffer from inefficient radio resource utilization, while fusion methods entail high implementation complexity and/or complex optimization problems. For communication-centric waveforms, emerging technologies like DDAM, OTFS, OCDM, and AFDM exhibit superior sensing performance in high-mobility scenarios but are often accompanied by complex signal processing and high computational demands.
Consequently, OFDM remains the most likely waveform for future ISAC wireless networks. In the 3GPP proposals [99, 100], an agreement was reached that the CP-OFDM waveform as defined in the 5G new radio (NR) is supported as the basis for the 6G radio (6GR) downlink, the CP-OFDM and DFT-s-OFDM waveforms as defined in 5G NR are supported as the basis for the 6GR uplink. For MIMO-OFDM ISAC systems, the use of XL-MIMO and large time-frequency bandwidths significantly enhances the sensing performance, while simultaneously introducing increased parameter estimation complexity. This challenge becomes particularly pronounced when performing joint estimation across the spatial, delay, and Doppler domains, compounded by the additional complexity of near-field parameter estimation. Thus, in this paper, we present a tutorial overview of rigorous and unified models for MIMO-OFDM ISAC, methods and corresponding algorithms for parameter estimation across the three domains, and analyses of their performance and complexity. Furthermore, the extension of far-field methods to near-field scenarios will be elaborated.
@@ III Fundamental ISAC Models
Considering the universal applicability of ISAC frameworks to all possible waveforms, we first introduce various ISAC modes, characterizing the relationships among transmitter and receiver signals and considering the required signal processing. As shown in Fig. 2, 3GPP standardization bodies have discussed six basic ISAC modes [101], which can be divided into two categories depending on whether the sensing signal is received and processed at the BS or UE side. For BS-side sensing, as shown in Fig. 2 (a)-(c), ISAC signals could be sent by either the BS itself, a UE, or another BS, corresponding to the BS monostatic, UE-to-BS bistatic, and BS-to-BS bistatic modes, respectively. Similarly, for UE-side sensing, as shown in Fig. 2 (d)-(f), ISAC signals could be sent by either the UE itself, a BS, or another UE, corresponding to the UE monostatic, BS-to-UE bistatic, and UE-to-UE bistatic modes, respectively. In the following, we present the signal input-output relationships for C&S based on a unified ISAC model that is applicable to all six modes above.
@@ III-A Unified ISAC Model
Fig. 3shows a generic ISAC setup that includes the aforementioned six modes as special cases. It consists of three components, namely an ISAC transmitter (ISAC-TX), a communication receiver (Com-RX), and a sensing receiver (Sen-RX). The ISAC-TX transmits ISAC signals to the Com-RX to provide communication services, and to the Sen-RX for sensing $K$ targets whose locations are denoted as $\mathbf{q}_{k},k=1,2,\cdots,K$. Denote the number of antennas at the ISAC-TX, Com-RX, and Sen-RX by $M_{t}$, $M_{c}$, and $M_{r}$, respectively. For sensing, we assume that each target has a LoS path with the ISAC-TX and the Sen-RX, while such an assumption is not needed for communication. The array response vector between the ISAC-TX and target $k$ is denoted as $\mathbf{a}_{\mathrm{T}}(\mathbf{q}_{k})\in\mathbb{C}^{M_{t}\times 1}$, which in general is a function of the sensing target location $\mathbf{q}_{k}$. Similarly, the RX array response vector is denoted as $\mathbf{a}_{\mathrm{R}}(\mathbf{q}_{k})\in\mathbb{C}^{M_{r}\times 1}$. Specific expressions for $\mathbf{a}_{\mathrm{T}}(\mathbf{q}_{k})$ and $\mathbf{a}_{\mathrm{R}}(\mathbf{q}_{k})$ in far-field and near-field scenarios will be provided in Section III-B.
@@
where $\ast$ denotes convolution, and $\mathbf{n}\bigl(t\bigr)\sim CN\bigl(0,\mathbf{\sigma}^{2}\mathbf{I}_{M_{c}}\bigr)$ is circularly symmetric additive white Gaussian noise (AWGN) with zero mean and variance $\sigma^{2}$. We assume communication signal processing methods consistent with traditional communication systems, and thus such methods will not be detailed here.
For target sensing, the ISAC signal sent by the ISAC-TX is reflected/scattered by the $K$ targets and other scatterers in the surrounding environment, and then received by the Sen-RX. A LoS link between the ISAC-TX and Sen-RX may also exist. Thus, the sensing signal $\mathbf{y}_{s}\left(t\right)\in\mathbb{C}^{M_{r}\times 1}$ received by the $M_{r}$-antenna Sen-RX array can be expressed as
@@
$\times e^{j2\pi\left(f_{\Delta}+\upsilon_{k}\right)t}+\mathbf{H}_{\mathrm{Clutter}}\left(t,\tau\right)\ast\mathbf{x}\left(t\right)+\tilde{\mathbf{z}}\left(t\right),$
where $\alpha_{k}$, $\tau_{k}$, and $\upsilon_{k}$ denote the complex-valued channel coefficient, the propagation delay, and Doppler frequency of the path corresponding to the $k$-th sensing target, respectively, $\tau_{\Delta}$ and $f_{\Delta}$ are the symbol time offset (STO) and carrier frequency offset (CFO) caused by the difference between the reference clock and local oscillators (LO) at the ISAC-TX and Sen-RX. Note that we may assume $\tau_{\Delta}=0$ and $f_{\Delta}=0$ for monostatic ISAC modes, while $\tau_{\Delta}\neq 0$ and $f_{\Delta}\neq 0$ in general for bistatic ISAC modes. The term $\mathbf{H}_{\mathrm{Clutter}}\left(t,\tau\right)$ is the channel impulse response of the clutter, which consists of the signals scattered/reflected by objects irrelevant to the $K$ sensing targets and the LoS link between the ISAC-TX and Sen-RX. Note that for monostatic ISAC systems, the ISAC-TX and Sen-RX are collocated and operate in full-duplex mode to achieve maximum SE, making them vulnerable to self-interference (SI) due to imperfect isolation. The residual SI after the application of various cancellation techniques can also be considered as a component of the clutter. Finally, $\tilde{\mathbf{z}}\left(t\right)$ is the AWGN at the Sen-RX.
@@
TABLE IV: The unified ISAC model and its special cases corresponding to the six basic modes.
@@
Note that except for some differences in the number of antennas and computational capabilities, BS-side and UE-side sensing are symmetric in terms of signal modeling and processing. Therefore, in the following, we present detailed models and signal processing techniques for ISAC systems of the three BS-side sensing modes based on the unified ISAC model, while similar models can be derived for UE-side sensing as well. To this end, we first show that the generic ISAC setup in Fig. 3 and the resulting signal in (3) can be reduced to the three BS-side sensing modes shown in Figs. 2 (a)-(c).
@@
  • BS monostatic ISAC mode: For BS monostatic ISAC mode, the ISAC-TX and Sen-RX shown in Fig. 3 are collocated at the BS. To maximize the SE, the ISAC BS needs to employ in-band full-duplex operation for simultaneous downlink ISAC signal transmission and uplink sensing signal reception. Note that while collocated at the BS, we assume that the ISAC-TX and Sen-RX use different sets of antennas for ease of SI cancellation. The Com-RX in Fig. 3 corresponds to a UE. In this case, the generic ISAC setup in Fig. 3 reduces to the BS monostatic setup in Fig. 4(a). Furthermore, the ISAC signal $\mathbf{x}(t)$ in (1)-(3) is transmitted by the BS, and is denoted as $\mathbf{x}_{\mathrm{BS}}(t)\in\mathbb{C}^{M_{t}\times 1}$.